OpenStudy (yanasidlinskiy):

The floor of a shed has an area of 99 square feet. The floor is in the shape of a rectangle whose length is 7 feet less than twice the width. Find the length and the width of the floor of the shed. Use the formula, area=length*width

OpenStudy (yanasidlinskiy):

length is 7 feet less than twice the width = 7-2x? Something like that...

OpenStudy (mathstudent55):

width = w twice the width: 2w 7 ft less than twice the width: 2w - 7

OpenStudy (mathstudent55):

A = LW = w(2w - 7) = 99

OpenStudy (yanasidlinskiy):

WAIT! I want to solve it lol. I need to understand it on my own on how to do it haha. Wait, why would it matter if I put it the way you showed me?

OpenStudy (mathstudent55):

There is a difference in the way you write it. Let me explain. Let's say Mary is 16 years old and Jane is 18 years old. You can say Mary's age is 2 years less than Jane's age. If that means 18 - 2, then Mary is 16 years old which is correct. If mistakenly you think that means 2 - 18 = -16, then Mary is -16 year old which is meaningless.

OpenStudy (mathstudent55):

"7 less than 2w" is 2w - 7 "7 less" means subtract 7, take away 7, from 2w 7 - 2w is "2w less than 7"

OpenStudy (yanasidlinskiy):

Gotcha! Ok, makes sense. Lemme solve this real quick. Sorry if it'll take a while..

OpenStudy (mathstudent55):

No problem. Take your time.

OpenStudy (yanasidlinskiy):

I got 22.

OpenStudy (mathstudent55):

w(2w - 7) = 99 2w^2 - 7w - 99 = 0

OpenStudy (yanasidlinskiy):

That's what I had in the first place..

OpenStudy (mathstudent55):

\(\large x = \dfrac{-b \pm \sqrt {b^2 - 4ac}}{2a} \) \(\large w = \dfrac{7 \pm \sqrt {(-7)^2 - 4(2)(-99)}}{2(2)} \) \(\large w = \dfrac{7 \pm \sqrt {49 + 792}}{4} \) \(\large w = \dfrac{7 \pm \sqrt {841}}{4} \) \(\large w = \dfrac{7 \pm 29}{4} \) \(\large w = \dfrac{36}{4} \) or \(w = \dfrac{-22}{4}\) \(\large w = 9\) or \(\large w = -5.5\)

OpenStudy (mathstudent55):

Since a width cannot be negative, we discard w = -5.5. The width is 9 ft. The length is 2w - 7 = 2(9) - 7 = 18 - 7 = 11 ft

OpenStudy (yanasidlinskiy):

It's the minor stuff that kills me lol.

OpenStudy (mathstudent55):

You need to work carefully. I went through the same, making stupid mistakes. Do your work and check it.

OpenStudy (mathstudent55):

When you find an answer, check the answer. We get W = 9 ft and L = 11 ft Let's see if it makes sense. The area has to be 99 ft^2 A = LW = 11 ft * 9 ft = 99 ft^2 The length has to be 7 less than twice the width. L = 2w - 7 = 2(9) - 7 = 18 - 7 = 11 ft Yes, L = 11 ft and W = 9 ft make sense.

OpenStudy (yanasidlinskiy):

Gotcha! Ok, makes sense!

OpenStudy (yanasidlinskiy):

Thanks for the detailed explanations. I appreciate it!

OpenStudy (mathstudent55):

You are very welcome.